OpenStudy (anonymous):

4. What is the simplest form of the expression?

OpenStudy (anonymous):

what expression

OpenStudy (anonymous):

\[\frac{ \sqrt{3}-{\sqrt{6}} }{ \sqrt{3} +{\sqrt{6}}}\]

OpenStudy (anonymous):

sorry idk bye

OpenStudy (anonymous):

A.\[\frac{ -1 - 2\sqrt{18} }{ 3 }\] B.\[\frac{ -3 - 2 \sqrt{18} }{ 9 }\] C,\[-3+2\sqrt{2}\] D.\[9-2\sqrt{18}\]

OpenStudy (anonymous):

\[\frac{ \sqrt{3}-{\sqrt{6}} }{ \sqrt{3} +{\sqrt{6}}}\times \frac{ \sqrt{3}-{\sqrt{6}} }{ \sqrt{3} -{\sqrt{6}}}\] is a start

OpenStudy (anonymous):

the denominator is \(3-6=-3\) and the numerator is \[(\sqrt3-\sqrt6)(\sqrt3-\sqrt6)\]\[=3-2\sqrt{18}+6\] \[=9-2\sqrt{18}\] \[=9-6\sqrt{2}\]

OpenStudy (anonymous):

final answer is got by reducing \[\frac{9-6\sqrt2}{-3}\]

OpenStudy (anonymous):

thats not one of my answer choises

OpenStudy (anonymous):

the answer is correct - just need to be simplified to get your choice. \[\frac{ 9 }{ -3 } + \frac{ -6\sqrt{2} }{-3 }\] = \[-3 + 2\sqrt{2}\] which is choice C

OpenStudy (anonymous):

thank you! can you help me with some more? i have about 18 more i have to do if you can help

OpenStudy (anonymous):

@VAlb